### v2

Abstract
In an era of increasing computational capabilities and growing environmental consciousness, organizations face a critical challenge in balancing the accuracy of forecasting models with computational efficiency and sustainability. Global forecasting models, lowering the computational time, have gained significant attention over the years. However, the common practice of retraining these models with new observations raises important questions about the costs of forecasting. Using ten different machine learning and deep learning models, we analyzed various retraining scenarios, ranging from continuous updates to no retraining at all, across two large retail datasets. We showed that less frequent retraining strategies maintain the forecast accuracy while reducing the computational costs, providing a more sustainable approach to large-scale forecasting. We also found that machine learning models are a marginally better choice to reduce the costs of forecasting when coupled with less frequent model retraining strategies as the frequency of the data increases. Our findings challenge the conventional belief that frequent retraining is essential for maintaining forecasting accuracy. Instead, periodic retraining offers a good balance between predictive performance and efficiency, both in the case of point and probabilistic forecasting. These insights provide actionable guidelines for organizations seeking to optimize forecasting pipelines while reducing costs and energy consumption.
keywords:
Time series, Demand forecasting, Forecasting competitions, Cross-learning, Global models, Machine learning, Deep learning, Green AI, Conformal predictions
1 Introduction
Forecasting plays a critical role in decision-making processes across industries, from supply chain management to energy demand planning. Traditionally, time series forecasting has relied on a Local Modeling approach (LM), meaning that separate models 111 We use the words model and method interchangeably, even if the two have distinct statistical meanings (Svetunkov & Boylan, 2023). are trained for each series in isolation. However, recent advances have demonstrated the potential of Global forecasting Models (GM), a paradigm also known as cross-learning, which in turn consists of fitting a single forecasting model on the whole data set (Januschowski et al., 2020). This novel approach allows to leverage information across multiple time series to improve accuracy and generalization by employing machine learning or deep learning architectures 222 We distinguish between traditional machine learning models, like linear regression or tree-based methods, and deep learning models, those that are based on some sort of neural network architecture. . Global modeling has affirmed itself as one of the most relevant innovations in time series forecasting of recent years (Semenoglou et al., 2021), and its success is not only due to the demonstrated forecast accuracy in most time series competitions, but also to the particular computational benefit it provides. Indeed, while the local approach to time series forecasting is potentially beneficial for forecast accuracy, it comes with substantial computational overhead. Moreover, since the forecasts are usually produced through some cloud computing service that relies on pay-as-you-go pricing (Fotios Petropoulos & Spiliotis, 2024), higher computing time and resources directly translate into higher costs of forecasting for organizations.
The retail industry particularly exemplifies the complexity of modern forecasting challenges. Retailers must predict demand for thousands or even millions of products across hundreds of stores while accounting for seasonality, promotions, pricing strategies, and external factors such as weather and local events. Accurate demand forecasts directly impact inventory decisions: overestimation leads to excess inventory and waste, particularly critical for perishable goods, while underestimation results in stockouts and direct money loss for companies (Fildes et al., 2022). Traditional approaches involving separate models for each product-store combination have proven increasingly inadequate, especially in contexts where the number of series (SKUs) to forecast is high, leading to the adoption of global forecasting models that can learn patterns across the entire product hierarchy and can drastically reduce the costs of producing forecasts 333 Most of the cost of producing forecasts is due to the algorithm’s training time. The global modeling approach cuts it by training the model only once on the whole data set. .
Nevertheless, despite the growing adoption of global models, it is still common practice to update the forecasting models when new observations are available, often motivated by the assumption that continuous updates lead to better adaptability to changing patterns and more accurate predictions. Unlike local models, global models benefit from learning shared dynamics across multiple time series, potentially leading to more robust representations that are less sensitive to frequent retraining. In the context of local models, the seminal work of Spiliotis & Petropoulos (2024) demonstrated that less frequent model updates do not harm forecasting accuracy. However, in the context of global models, the effects of retraining are not well understood. Is continuous retraining necessary to maintain forecast performance, or can global models remain effective without frequent updates? This aspect remains largely unexplored in the forecasting literature, and investigating the impact of retraining is crucial for both methodological and practical reasons. From a methodological perspective, understanding whether global models degrade in performance without frequent retraining provides insights into their stability and adaptability. If global models can maintain strong predictive accuracy with less frequent updates, it would challenge conventional wisdom on the necessity of continuous retraining in forecasting. Moreover, a stable forecasting model provides great benefit to the whole decision process the forecasts are used for. From a practical perspective, frequent retraining has significant computational and environmental costs. Training large-scale forecasting models requires substantial computational power, contributing to energy consumption and carbon emissions (Schwartz et al., 2020). Indeed, the energy consumption of model training extends beyond direct computational costs. In recent years, the environmental impact of machine learning models has become a growing concern. Data centers running these models contribute significantly to global carbon emissions, with estimates suggesting that training a single large deep learning model can emit as much carbon as five cars over their lifetimes (Strubell et al., 2019). The frequency of model retraining multiplies this impact, particularly in large organizations handling millions of series, emphasizing the need for more computationally efficient training practices. The implication is straightforward: reducing the retraining frequency of forecasting models could contribute substantially to sustainability by lowering energy consumption, possibly without harming accuracy performance.
1.1 Research Question
We aim to extend the results presented in the influential work of Spiliotis & Petropoulos (2024), addressing the question ”Is frequent retraining necessary in the context of global forecasting models?”. Specifically, we study whether skipping retraining when new observations are available harms the forecasting performance of global models. To answer this question, we use the two most recent and comprehensive retail forecasting datasets, namely the M5 and the VN1 competitions’ data.
To generally understand how retraining the forecasting models affects their performance, we consider ten different global forecasting methods (five from the ”classical” machine learning domain and five often-used deep neural network architectures), and several possible retraining scenarios, from continuous retraining to no retraining at all. We also explore intermediate periodic retraining strategies to broadly cover the most reasonable and effective scenarios.
We also focus on the investigation of trade-offs between accuracy and sustainability, in terms of the computational cost of resources to produce the forecasts. This cost is indeed significant for large-scale applications, like the retail industry, and being able to reduce (or control) it somehow may result in direct and significant business savings.
1.2 Contributions
Our contribution is threefold:
  • 1.
    We provide the first comprehensive study of the relationship between retraining frequency and forecast accuracy using 10 different global models, a diverse set of real-world datasets, and focusing on both point and probabilistic forecasting.
  • 2.
    We compare different retraining scenarios (e.g., continuous, periodic, and no retraining) on different datasets to quantify the impact of frequent retraining in terms of the cost of forecasting.
  • 3.
    We present practical guidelines for organizations and practitioners on when and how often to retrain global forecasting models to balance accuracy and cost.
By addressing these points, this paper contributes to both the forecasting and machine learning communities, offering insights into the trade-offs between accuracy, efficiency, and sustainability in global forecasting models.
1.3 Overview
The rest of this paper is organized as follows. After a brief review of related works (Section 2), in Section 3, we describe the design of the experiment used in our study. The datasets and their characteristics are presented in 3.1, the methods adopted for global forecasting are discussed in 3.2, and the concepts related to model update and retrain scenario are explained in 3.3. The evaluation setup through rolling origin validation is presented in 3.4, and the metrics used to evaluate the model performances are shown in 3.5. In Section 4, we discuss the empirical findings of our study, including forecast accuracy, computing time, and cost analysis of the different scenarios. Finally, Section 5 contains our summary and conclusions.
2 Related works
The literature on cross-learning approach has evolved significantly in recent years. Nowadays, most of the works related to time series forecasting include at least some benchmark comparison with global models, demonstrating their relevance in the field. Semenoglou et al. (2021) extensively showed their accuracy on the M4 competition dataset, Hewamalage et al. (2022) evaluated the conditions when global forecasting models are competitive, and Montero-Manso & Hyndman (2021) and Montero-Manso (2023) theoretically demonstrated that GM are at least as accurate as local models with less complexity and without any assumption on the similarity of the data. GM emerged as the most accurate approach in many forecasting areas, such as gas consumption (Gaweł & Paliński, 2024), electricity demand (Buonanno et al., 2022), water demand (Groß & Hans, 2024), crop production (Ibañez & Monterola, 2023), and retail demand (Spiliotis et al. (2022), Bandara et al. (2019), Juan R Trapero & Fildes (2015)). Moreover, GM effectiveness shined during the M5 competition (Makridakis et al., 2022a), where tree-based models leveraging cross-learning were among almost all the top forecasting solutions (Januschowski et al., 2022). Several techniques like clustering (Godahewa et al. (2021a), Bandara et al. (2020)), and data augmentation (Bandara et al., 2021) have been tested to further increase the performance of GM. Furthermore, new machine learning (Godahewa et al., 2023) and deep learning (Oreshkin et al., 2020) architectures specifically designed for cross-learning have been developed. Finally, a novel area of research is emerging that tries to improve the ability of GM to capture local patterns (Sen et al., 2019), and their explainability (Rajapaksha et al., 2023).
From a forecasting evaluation perspective, most of the literature on GM is focused on point prediction accuracy, possibly because most machine learning and deep learning methods do not directly output probabilistic forecasts (Makridakis et al., 2022c). Nevertheless, in many forecasting contexts (like supply chain) it is very important to be able to produce and evaluate predictions in a probabilistic way (being intervals, quantiles, or density based) (Fildes et al., 2022). Among others, Vovk et al. (2005) introduced a novel tool for uncertainty quantification under any machine learning model, namely Conformal Inference, that can also be applied in time series forecasting experiments (Stankeviciute et al., 2021).
In the context of model retraining and updating strategies, instead, Spiliotis & Petropoulos (2024) is the main work related to time series forecasting. The authors explored extensively the effects of different retraining scenarios and different forms of model parameter updates on the model performance, although they focused on the exponential smoothing family of models following the traditional local approach. Huber & Stuckenschmidt (2020) briefly discussed retraining in the context of retail demand, but with few models and retraining scenarios, and on a proprietary daily dataset only. Despite the findings being promising, there has been little direct exploration of whether global models specifically require frequent retraining or if they can retain competitive accuracy with less frequent updates. However, this topic is under consideration of the broader machine learning community (Getzner et al., 2023), advocating for Green AI (Schwartz et al., 2020).
Our study builds upon these existing works, directly investigating the necessity of retraining in global forecasting models. By evaluating many different retraining strategies and their impact on the forecast accuracy of several global models, we aim to provide both theoretical insights and practical recommendations for more sustainable forecasting practices.
3 Experimental design
This section describes the empirical analysis we conducted to study whether less frequent retraining scenarios may produce similar accuracy results compared to the baseline scenario (that with the highest retraining frequency). First, we describe the datasets used in the experiments, and then explain the different machine learning and deep learning models adopted. The performance measures, the several possible scenarios, and the strategy used to evaluate the forecasts are also discussed.
3.1 Datasets
For the experiment, we used two retail forecasting datasets: the M5 and the VN1 competition datasets. The M5 competition was organized by Spyros Makridakis and his colleagues as part of the M-competitions series, which aimed to compare different forecasting methods in the context of retail demand (Makridakis et al., 2022b). The M5 (Howard & Makridakis, 2020) is a well-known and well-studied dataset containing 30,490 SKU $\times$ Store daily time series of Walmart’s unit sales of products. It covers the sales of three categories of products (Food, Hobbies, and Household) sold into ten different stores located in three US states (California, Texas, and Wisconsin). The time period spans from 2011 to 2016. The time series are highly intermittent and are hierarchically organized, allowing forecasting at multiple levels, such as individual products, product categories, stores, and States. Exogenous information that can influence sales, such as product prices, promotions, and special events (e.g., holidays) are also available. The VN1 Forecasting - Accuracy Challenge competition was jointly organized by Flieber, Syrup Tech, and SupChains in October 2024, and it is the first of its series (Vandeput, 2024). The dataset contains weekly sales of 15,053 products sold from 2020 until 2024 from e-vendors. In particular, they were mostly online from the US, and the products were directly sold to the final consumers. Contrary to the M5 dataset, where all the products were sold by a single retailer (Walmart) and from just a few stores, the VN1 dataset includes product sales of 328 warehouses from 46 different retailers. As far as we know, we are among the first to test forecasting models on this data. These sets of data are the most recent and comprehensive time series datasets related to retail demand, allowing for a good generalization of the results, in particular in the context of demand forecasting.
Table 1: Characteristics of the different datasets used in the experiments. The number of series is based on the minimum number of observations per series.
Dataset Frequency N. Series Min Obs per Series
M5 Daily (7) 28.298 730
VN1 Weekly (52) 15.053 157
In both cases, in our experiment, we focused on the most disaggregated level (SKU $\times$ Store), since the potential benefit of retraining the models less frequently is much larger at lower levels of aggregation. Moreover, for the M5 competition dataset, we did not consider the whole set of time series to be able to consistently apply the evaluation setup described in Sections 3.4. In particular, for daily data, we kept only time series with at least two years of data (730 observations), while for weekly data we considered only those that had at least three years of data (157 observations).
3.2 Forecasting models
In this section, we provide an overview of the global models employed for our experiments.
Let us define $\mathcal{Y}$ as the set of all available time series in a dataset, such that $Y_{i}$ represents a single component, and let $\mathcal{F}$ be the set of possible predictive functions, so that $F$ corresponds to a single model 444 In our setting, $F$ can be any predictive model in machine learning or deep learning framework. We do not consider classical statistical forecasting methods, like ARIMA or Exponential Smoothing, since in their common formulations are purely local models. . Without loss of generality, we can assume that all the necessary information for prediction are contained in $\mathcal{Y}$. Under the local approach, predictions for the forecast horizon $h$ are obtained training a model for each time series in the data set, implying that each time series has its own model, defined by its own parameters values.
$Y_{i}^{h}=F(Y_{i},\theta_{i})\mbox{.}$ (1)
On the contrary, following the global modeling framework, forecasts for each time series are produced by a model trained on the whole data set.
$Y_{i}^{h}=F(\mathcal{Y},\Theta)\mbox{.}$ (2)
Notice how in the cross-learning methodology, the parameters $\Theta$ are not series-specific, but are common to all time series.
In our experiment, we focused on analyzing the performance of global methods only since we are interested in testing whether this approach, as opposed to the local one, can benefit from retraining the models less frequently. Indeed, nowadays, cross-learning is the go-to approach for almost all the industries involved with huge time series datasets, as in the context of retail demand forecasting, where the forecasts of many different SKUs have to be provided regularly (Januschowski et al., 2020). For a comprehensive evaluation of global modeling approaches in demand forecasting, we used both traditional machine learning models and cutting-edge deep learning methodologies. The models were selected for their established performance in time series forecasting tasks and their diverse methodological approaches, enabling a wide comparison and avoiding results being dependent on a specific algorithm. Furthermore, as benchmarking against established methods is essential in time series forecasting, we included two widely recognized global forecasting models, Linear (Pooled) Regression and Multi-Layer Perceptron, as baselines to evaluate the performance of the proposed approaches.
All the global models were trained using Python under Nixtla’s framework (Nixtla, 2022). The mlforecast and the neuralforecast libraries were used to train machine learning and deep learning models efficiently.
3.2.1 Machine learning models
Machine learning models have demonstrated their effectiveness in forecasting tasks due to their ability to capture non-linear relationships in the data. Moreover, they are often easy to train and usually produce very accurate results, leveraging the cross-learning approach. In this study, we experimented with Linear (Pooled) Regression and four different tree-based methods.
Linear Regression (LR) is a classical statistical model that assumes a linear relationship between input features and the target variable. Despite its simplicity, LR can be effective for time series forecasting when combined with appropriate feature engineering. The Pooled Regression is considered a solid benchmark for global model performance evaluation (Montero-Manso & Hyndman, 2021; Godahewa et al., 2021b), and it has also proven to be quite effective (Bandara et al., 2022).
Random Forest (RF) is an ensemble learning method based on regression trees. It constructs multiple trees during training and aggregates their predictions through averaging. RF excels in capturing non-linear patterns and interactions between variables, and it is robust to overfitting, making it a strong tree-based model for demand forecasting (Breiman, 2001). It was the method used by Amazon until 2015 to forecast e-commerce products demand (Januschowski et al., 2022) 555 For the cost of computation, the Random Forest model is tested only on the VN1 weekly dataset. .
Extreme Gradient Boosting (XGBoost) is a gradient-boosted decision tree algorithm designed for speed and performance. Iteratively optimizing an objective function combines weak learners’ predictions to form a solid predictive model (Chen & Guestrin, 2016). XGBoost incorporates regularization techniques to prevent overfitting, making it particularly effective for datasets with complex relationships (Shwartz-Ziv & Armon, 2022).
Light Gradient Boosting Machine (LGBM) is another gradient-boosted tree algorithm that emphasizes efficiency and scalability. It uses a histogram-based approach to split features and a leaf-wise growth strategy to reduce computation time. LGBM is particularly well-suited for large datasets and high-dimensional data, often outperforming other boosting algorithms in terms of speed and accuracy (Ke et al., 2017). In demand forecasting, it is among the top solutions of all the major competitions (Favorita, Rossmann, M5, and VN1) (Makridakis et al., 2022a, c; In & Jung, 2022; Lainder & Wolfinger, 2022).
Categorical Boosting (CatBoost) is a gradient-boosted decision tree algorithm specifically designed to handle categorical data. By leveraging ordered boosting and other innovations, it fosters accuracy and enhances generalization (Prokhorenkova et al., 2018). CatBoost’s ability to handle categorical features without extensive preprocessing makes it advantageous for forecasting tasks involving categorical covariates. During the last years, it consistently showed top performance on many tabular data studies, becoming the go-to solution for many practitioners in the field (Shmuel et al., 2024; Ye et al., 2025).
Machine learning models have the advantage of being easier to train with respect to their deep learning alternatives. However, they require extensive and careful feature engineering to produce accurate results (Januschowski et al., 2022). For this reason, we followed simplified versions of the M5 and VN1 top solutions to build time series features. In particular, we used lags, rolling averages, expanding averages, calendar features (year, quarter, month, week, day of week, day), and static features (store, category, location, product identifiers) based on the frequency of the dataset. Moreover, for the M5 data, we also used external features related to special events as available. The most relevant hyperparameters of the models were selected based on top-performing solutions, otherwise the default values suggested by the software providers were adopted.
3.2.2 Deep learning models
Deep learning models have gained prominence in time series forecasting due to their capacity to model longer-term dependencies in the data and to easily learn hierarchical representations from raw data (Goodfellow et al., 2016). In our experiment, we employed five different neural network architectures, two well-known methods and three state-of-the-art models.
Multi-Layer Perceptron (MLP) is a feedforward neural network consisting of an input layer, one or more hidden layers, and an output layer. Each layer applies a non-linear activation function to capture complex relationships in the data (Rosenblatt, 1958). MLPs are a versatile and very efficient solution. For this reason, many deep learning models specifically created for time series forecasting are MLP-based. MLP is also considered a solid benchmark for deep learning global forecasting models (Godahewa et al., 2021b).
Recurrent Neural Networks (RNN) are designed to model sequential data by maintaining a hidden state that captures temporal dependencies. Standard RNNs, however, suffer from vanishing gradient problems, limiting their ability to learn long-term dependencies (Cho et al., 2014). Variants such as Long Short-Term Memory (LSTM) networks and Gated Recurrent Units (GRUs) address this limitation and are widely used for time series forecasting (Hochreiter & Schmidhuber, 1997) 666 In our work, we used only the LSTM version, but, for the cost of computation, it was trained only on the VN1 weekly dataset. .
For a long time in deep learning, sequence modeling was synonymous with recurrent networks, yet several papers have shown that simple convolutional architectures can outperform canonical recurrent networks like LSTMs by demonstrating longer effective memory. Temporal Convolutional Networks (TCN) are convolutional architectures tailored for sequential data that, by employing causal convolutions and dilations, capture long-range dependencies (Van den Oord et al., 2016).
Neural Basis Expansion Analysis for Time Series (NBEATS) is a deep learning model specifically designed for time series forecasting. It employs a sequence of fully connected layers organized into blocks. Each block outputs both trend and seasonal components, enabling the model to decompose and predict time series effectively. The network has an interpretable configuration that sequentially projects the signal into polynomials and harmonic basis to learn trend and seasonality components (Oreshkin et al., 2020). The Neural Basis Expansion Analysis with Exogenous (NBEATSx), allows for the inclusion of exogenous temporal variables available at the time of the prediction (Olivares et al., 2023). This method has shown state-of-the-art performance on several benchmark datasets and competitions (Anderer & Li, 2022).
Neural Hierarchical Interpolation for Time Series (NHITS) builds upon the success of NBEATS by incorporating hierarchical interpolation mechanisms to better capture time hierarchies in time series data. Multi-rate input pooling, hierarchical interpolation, and backcast residual connections together induce the specialization of the additive predictions in different signal bands, reducing memory and computational time, thus improving the architecture’s parsimony and accuracy (Challu et al., 2022).
Deep learning models typically do not require extensive feature engineering step needed by their machine learning counterpart, since they create such features (like lags and rolling averages) internally. Nevertheless, they usually are more difficult to train since they have many more hyperparameters to select affecting the forecasting performance (Smyl, 2020). We trained the global deep learning models by adding static, calendar, and external features only, and relying on the top competitions’ solutions to set the hyperparameters’ values.
3.3 Retrain scenarios
To answer our research question, we explored several possible retraining scenarios. A retrain scenario or retrain window, $r$, is identified as a positive integer representing the frequency at which the model is re-trained or updated. Essentially, it represents how many data points need to be passed before a new training step is performed. The retrain scenarios strongly depend on the frequency of the dataset considered because the frequency drives both the forecast horizon and the business review periods. Therefore, we defined different set of retrain scenarios based on the frequencies of our datasets:
  • 1.
    for weekly data, $r=\{1,2,3,4,6,8,10,13,26,52\}$
  • 2.
    for daily data, $r=\{7,14,21,30,60,90,120,150,180,364\}$
For instance, in the case of daily data, $r=7$ implies that the model is retrained every 7 new observations come in, that is, every week. Each set contains ten different values, being as exhaustive and computationally feasible as possible.
Note that we are training global models, hence the training dataset is composed by the training set of each time series in the dataset. Since the datasets we considered are all aligned, which implies that, at every retrain period, we fit the model on a new training data set containing $r$ new observations for each time series (where $r$ is the chosen retrain scenario). Moreover, note also that we are considering only two different forms of model update: a model can either be completely updated when retraining is performed, or used as is to produce forecasts for $r$ subsequent periods. Therefore, we are not examining the effect of hyperparameter tuning within each retrain scenario, given the expensive computational cost of this process and minor changes expected.
The scenario $r=1$ is the so-called continuous retraining, and it is the most expensive since it implies that the model is retrained for every new observation. Theoretically, but usually also practically, it should be the most accurate forecasting scenario, because the model used to predict has been trained on all the available data points up to that moment. For this reason, we considered this scenario as the benchmark both in terms of forecasting accuracy and computational cost. For daily data only, however, the benchmark scenario is $r=7$, because it is not common to retrain a global model every day, and usually the update is performed once a week. The scenario $r=T$ is the no retraining scenario, meaning that the forecasting model is fitted just once on the initial training set, then it is used to produce the forecasts for the entire test set $T$. It is the lowest computationally expensive but it should also be the lowest accurate. All the other scenarios such that $1<r<T$ are considered as periodic updating strategies. Both the accuracy and the computing time should be non-increasing functions of $r$.
3.4 Evaluation setup
In time series forecasting, out-of-sample testing is crucial for assessing how well a model can generalize to unseen data. This is particularly important because patterns in historical data may not fully represent future trends, and unexpected changes in the data may occur (Tashman, 2000). As suggested by Bergmeir & Benítez (2012), rolling origin evaluation is the most widely used and correct method for conducting such tests, offering a systematic way to evaluate forecasting models over multiple iterations. The rolling origin evaluation process begins by dividing the time series into two parts: a training set, which is used to build the model, and a test set, which is used for evaluation. The test set always follows the training set chronologically to maintain the temporal order of the data. At each step, the model is trained on the training set and generates forecasts for a specified number of future points, the forecast horizon $h$. These forecasts are then compared to the actual values in the test set using a chosen evaluation metric (see Section 3.5). What makes rolling origin evaluation unique is its iterative nature. After each forecasting step, the forecast origin (the last point in the training set) is shifted forward by a fixed number of periods, the step size. The model is then retrained using the updated training set, which may either expand to include all available data up to the new forecast origin or remain fixed to a specific window size. This process is repeated until the test set is fully utilized, and the overall performance of the model is calculated by averaging the results across all the iterations.
Compared to the fixed origin evaluation setup, which allows for a single evaluation step, the rolling origin evaluation is preferable for its ability to provide a more comprehensive assessment of a model’s performance under different conditions, such as varying seasonal patterns, level shifts, or data trends. By simulating multiple forecast cycles, indeed, this method reduces the risk of bias that can arise from evaluating a model a single time and at a single forecast origin (Bergmeir & Benítez, 2012). This approach is beneficial in industries like retail and supply chain management, where forecasts need to be continuously updated to reflect real-time changes. By simulating multiple scenarios and evaluating a model’s performance across different cycles, rolling origin evaluation provides valuable insights into the robustness and reliability of the forecasting methods adopted.
In practice, the rolling origin evaluation setup can be adjusted based on the data and forecasting goals. For instance, the training set can be fixed to a rolling window of the most $n$ recent observations (fixed window), or it can expand to include the full historical data (expanding window) (Bergmeir & Benítez, 2012). In real-life applications, most forecast practitioners use this second option, especially when the length of the time series is small (Petropoulos & et al., 2022). In our study, we adopted the expanding window approach because we are interested in simulating a forecasting experiment as close as possible to the business reality. Moreover, it is also the only reasonable option when dealing with short time series, like the case for the weekly VN1 datasets.
Table 2: Retraining scenarios, test windows, and forecast horizons for each dataset. The setup is based on the respective time series frequency.
M5 VN1
Training Window ($w$) $\geq 730$ $\geq 157$
Test Window ($T$) 364 52
Horizon ($h$) 28 13
Step size ($p$) 1 1
Retraining Scenarios ($r$) {7, 14, 21, 30, 60, 90, 120, 150, 180, 364} {1, 2, 3, 4, 6, 8, 10, 13, 26, 52}
The size of the test set $T$, the length of the forecast horizon $h$, and the step size $p$, for shifting the forecast origin, are also customizable parameters that depend on the specific use case. Table 2 shows the parameters selected in our experiments. Following Spiliotis & Petropoulos (2024), we set the size of the test set to cover at least one complete year so that the evaluation of the different scenarios does not depend on intra-year variations, such as any specific season or period of the year. We selected the forecast horizons based on the type of business decisions that usually daily and weekly forecasts support, i.e. operational vs strategic planning. Finally, the step size was always set equal to one to maximize the number of evaluations of each scenario.
3.5 Performance metrics
The accuracy evaluation of point predictive models is a controversial topic in time series forecasting: many metrics are available to capture models’ performances, but no consensus has been reached in the literature on whether one metric is better than others (Hewamalage et al., 2023). However, since we are dealing with demand forecasting at SKU-level, implying that data intermittency is very much likely, all the metrics based on absolute or percentage errors are not optimal since they optimize for the median (Kolassa, 2020). Moreover, being the scale possibly different among the series, a scaled accuracy measure is preferable. For this reason, we considered the Root Mean Squared Scaled Error (RMSSE), proposed by Hyndman & Koehler (2006), to evaluate the point forecast accuracy of the models. The RMSSE is defined as:
$\text{RMSSE}=\sqrt{\frac{\frac{1}{h}\sum_{t=n+1}^{n+h}(y_{t}-\hat{y}_{t})^{2}}{\frac{1}{n-s}\sum_{t=s+1}^{n}(y_{t}-y_{t-s})^{2}}}\mbox{.}$ (3)
The RMSSE measures the relative prediction accuracy of a forecasting method by comparing the mean squared errors of the prediction and the observed value against the mean squared errors of the seasonal naive model. This metric was the official metric used to assess the performance of models in the M5 competition (with $s=1$, hence using the naive model as a benchmark) (Makridakis et al., 2022b). Lower RMSSE values suggest higher model accuracy.
Our work is not limited to the evaluation of point forecasting accuracy, but we also aim to assess the probabilistic performance of the models throughout each different retrain scenario. However, since the machine learning and deep learning models adopted for the experiments do not output density (or probabilistic) predictions by default, we relied on Conformal Inference as a general framework to build prediction intervals. Conformal Inference is a powerful tool that allows quantification of the uncertainty associated with out-of-sample predictions simply using a validation set on a point forecaster model to produce intervals at the desired levels (Vovk et al., 2005). It was originally proposed for non-temporal data, because the data exchangeability assumption, nevertheless, in recent years it has been extended to be used also in the context of time series data (Stankeviciute et al., 2021). In particular, given its guaranteed coverage, distribution-free, model-agnostic, computational efficiency, and low dataset size requirements, Conformal Inference is likely the only tool available to evaluate the uncertainty in a global forecasting setting where different models are compared among several datasets and scenarios. One may prefer, instead, to train models using quantile losses to directly yield quantile forecasts, but this would require training a single model for each quantile. Even if it may be a valuable way of modeling uncertainty, as demonstrated by the M5 Uncertainty competition (Makridakis et al., 2022c; Lainder & Wolfinger, 2022), it has the huge disadvantage of being very computationally expensive, hence negatively impacting on the sustainability of forecasting applications in large-scale environments. Moreover, even if multi-quantile losses can provide multiple quantiles, and prediction intervals, they usually are uncalibrated, meaning that the actual frequency of observations falling within the interval does not align with the confidence level associated with it. Therefore, for these multiple reasons, we decided to adopt Conformal Inference as general tool to estimate quantile forecasts.
Once the quantiles are obtained, we used a scaled version of the Multi-Quantile Loss to comprehensively measure the goodness of the probabilistic predictions. The scaled Quantile Loss, also known as (scaled) Pinball Loss, and the scaled Multi-Quantile Loss are defined as:
$\text{SQL}=\frac{1}{h}\frac{\sum_{t=n+1}^{n+h}\left(q\cdot(y_{t}-\hat{y}_{t})\cdot\mathbb{I}_{y_{t}\geq\hat{y}_{t}}+(1-q)\cdot(\hat{y}_{t}-y_{t})\cdot\mathbb{I}_{y_{t}<\hat{y}_{t}}\right)}{\frac{1}{n-s}\sum_{t=s+1}^{n}|y_{t}-y_{t-s}|}\mbox{,}$ (4)
$\text{SMQL}=\frac{1}{|\mathcal{Q}|}\sum_{q\in\mathcal{Q}}\text{SQL}(q)\mbox{,}$ (5)
where $\mathcal{Q}$ is the set of possible quantiles. The QL, being a proper scoring rule, allows for the correct evaluation of the probabilistic forecasts (Kolassa, 2016). The scaling factor is the in-sample, one-step-ahead mean absolute error of the seasonal Naive model. The scaled MQL (weighted and with $s=1$) was also the official metric used to evaluate the overall performance of the forecasting methods during the M5 Uncertainty competition (Makridakis et al., 2022c). For daily data we used $s=7$, because $s=1$ is arguably an irrelevant benchmark in daily retail data, which typically exhibits strong day-of-week seasonality (Fildes et al., 2022). Instead, we set $s=1$ for weekly data.
Note that we considered a total of 14 different quantiles, $q$, of probability levels {0.005, 0.025, 0.050, 0.100, 0.150, 0.200, 0.025, 0.750, 0.800, 0.850, 0.900, 0.950, 0.975, 0.995}. Therefore, we studied the 50% 60%, 70%, 80%, 90%, 95%, and 99% central prediction intervals. The 50% and 60% central prediction intervals provide a good description of the center of the forecast distribution, while the 90%, 95%, and 99% give more information about its tails, which in retail demand forecasting problems are fundamental to determine the appropriate safety stock levels (Barrow & Kourentzes, 2016). These quantiles provide sufficient information about the uncertainty of the forecasts and allow for the effective description of the whole distribution.
Note also that we computed conformal prediction intervals on a validation set at least twice as long as the forecast horizon. This was the main reason that forced us to reduce the number of time series in the datasets, that is to be able to obtain a reliable estimate of the model uncertainty through prediction intervals 777 We used a validation set four times larger than the forecast horizon for the M5 dataset and twice the forecast horizon for the VN1 dataset. This limited the number of time series from 30,490 to 28,298 for the M5 dataset. .
Finally, we addressed the problem of evaluating the cost associated to producing the forecasts by calculating the Computing Time (CT). CT is defined as the time in seconds required for training and predicting the next $h$ time steps ahead with the model. Among the possible measures of complexity, such as the number of parameters, number of iterations, model’s depth, etc, CT is the most directly related to the monetary cost of forecasting since the forecasts are usually produced through some sort of cloud computing service that relies on pay-as-you-go fares (Spiliotis & Petropoulos, 2024). Hence, CT is analyzed for each forecasting model and retraining scenario combination. Lower CT values imply lower forecasting costs.
Note that we related each evaluation metric result to the baseline retraining scenario, depending on the frequency of the dataset, to be able to easily compare the performance among different models and retrain windows. Furthermore, to statistically verify our research question, we tested the scenarios’ results using the Friedman-Nemenyi test for multiple comparisons (Demšar, 2006).
For the experiments, we used a cloud computing machine NC6s_v3 hosted on Microsoft Azure, with Linux Ubuntu 24 operating system, 1 Graphical Processing Unit, 6 Computing Processing Units, 112GB of memory. Parallelization and GPUs were used whenever possible within the model libraries. The utilsforecast library was used to evaluate the model performance.
4 Results and discussion
In general, the models tested performed better, in absolute terms, on the M5 dataset compared to the VN1 (see the RMSSE and SMQL tables in the Supplementary materials). This may be due to different factors like the dataset size, the frequency of the time series, the availability of external regressors to account for promotions, events, etc, and the absence of reference hyperparameter values to use during the training process.
Figure 1 shows the forecast accuracy of each model along the different retraining scenarios for the M5 and the VN1 datasets. To simplify comparisons, for each dataset, we report the results in relative terms with respect to the corresponding baseline scenario, that is $r=7$ for M5 and $r=1$ for VN1. With the exception of the CatBoost model, the RMSSE profiles are very stable over the retraining frequencies. Indeed, the accuracy remains practically the same in the M5 dataset and even improves in the VN1 dataset, regardless of the retraining scenario considered. Especially for low periodic retraining scenarios, the performance of most global models is indistinguishable from the baseline, and even if some deterioration is present for higher retraining frequencies, this is less than 5% also for the no retraining setup. These results imply that less frequent retraining does not harm the point forecast accuracy of global models. This can be explained by the fact that, if the demand patterns remain stable without significant trends or concept drifts, as is the case for both the M5 and VN1 datasets, the forecasts will accurately track demand over extended periods.
Figure 2 statistically confirms the above results for the M5 dataset 888 The Friedman-Nemenyi test is usually adopted to compare and rank the performance of different models. Here we use it to compare the accuracy produced by the same model over different retraining scenarios. . It is clear that the retraining scenarios are not statistically different in terms of point forecast accuracy at the 5% level. Indeed, even though the mean rank of the continuous retraining scenario is lower, the intervals intersect with the other scenarios, hence we can not differentiate between them. For the VN1 dataset (see Supplementary materials), less frequent retrainings are even statistically significant, meaning that periodic retraining does improve accuracy over continuous retraining.
Figure 1: RMSSE results for each method and retrain scenario combination in relative terms with respect to the baseline scenario, $r=7$ for the M5 dataset and $r=1$ for the VN1 dataset.
Figure 2: M5 Friedman-Nemenyi test results based on RMSSE.
In a similar fashion to Figure 1, Figure 3 summarizes the relative accuracy in a probabilistic forecasting setting, as defined in 3.5. In this context, we observe that, for the M5 dataset, the accuracy (as measured by Multi Quantile Loss) is clearly an increasing function of the retraining scenario, meaning that less frequent updates harm the probabilistic forecasting performance of the models. This is true regardless of the method used. Nonetheless, the differences in accuracy are irrelevant for low retraining scenarios and slightly more pronounced for higher retraining levels, but in any case, less than 5-6 percentage points. We also note a small difference in the performances of machine learning and deep learning models, where the formers perform consistently better as the retraining scenario increases. For the VN1 dataset, instead, we observe an almost convex relationship between the accuracy and the retraining period. On average, models’ performance improves for low retraining scenarios and then starts deteriorating around $r=4$. The only exceptions are NBEATSx and NHITS models, designed also for long-term forecasting, which are consistently better compared to the others for longer retraining scenarios, both for point and probabilistic forecasting.
The statistical test to compare probabilistic forecasting performances among the different retraining scenarios on the M5 dataset, as depicted in Figure 4, confirms that there are significant differences in accuracy. In particular, lower restraining scenarios tend to produce more accurate probabilistic forecasts, that is, less frequent retraining effectively harms the accuracy. However, this is not true for low retraining scenarios where the differences are not statistically significant, implying that even in the context of probabilistic forecasting the frequency of retraining may be reduced to at least once a month (or every two months). For the VN1 (results in Supplementary materials), instead, some level of periodic retaining is better than continuous retraining, most of the time.
Figure 3: SMQL results for each method and retrain scenario combination in relative terms with respect to the baseline scenario, $r=7$ for the M5 dataset and $r=1$ for the VN1 dataset.
Figure 4: M5 Friedman-Nemenyi test results based on SMQL.
Figure 5 shows the relative computational time of each retraining scenario for the two datasets. We observe that CT decreases exponentially as the retraining scenario increases. On average, reductions are similar for both, the M5 and the VN1: going from the baseline to the first periodic retraining scenario ($r=14$ or $r=2$) almost halves CT, retraining the models every month often reduces computing time by 75%, and this reduction reaches 90% in the no retraining scenario, where the models are trained just once. However, we observe that there is a strong difference between machine learning and deep learning models in the M5 data. While the former continuously undergoes CT reductions, the latter seems to plateau at 50%, experiencing much lower gains over increasing retraining scenarios. As expected, this difference is less evident in the VN1 dataset, since it is smaller and of lower frequency compared to the M5. These results will have direct implications in terms of the costs of forecasting, which in turn may guide the choice of the model to adopt.
Figure 5: CT results for each method and retrain scenario combination in relative terms with respect to the baseline scenario, $r=7$ for the M5 dataset and $r=1$ for the VN1 dataset.
Overall, the results of Figures 1, 3, and 5, combined with the Friedman-Nemenyi tests, suggest that retraining the global models less frequently does not harm (or even improves) forecast accuracy, while at the same time significantly reducing the computing time of producing the forecasts. Thus, extending the retraining time from the standard practice of continuous retaining to some level of periodic retraining allows for effective management of the computing time, and in turn, the costs of forecasting. Indeed, as already observed, the computational time can be directly translated into actual costs for the company. Following Nikolopoulos & Petropoulos (2018) and Fotios Petropoulos & Spiliotis (2024), we assumed some standard costs for computing services to estimate the costs of forecasting associated with each retrain scenario. Note that costs and savings are normalized by the number of series considered in each dataset (28,298 and 15,053) so that they can be directly compared and conclusions can be drawn in terms of the frequency of the time series. Figures 6, and 7 show the costs and associated savings for a large retailer, given a fixed computing service cost of \$3.5/hour, 200,000 unique SKUs and 5,000 stores, which is approximately the size of the forecasting problem of Walmart (Spiliotis & Petropoulos, 2024). As expected, the costs decrease exponentially with less frequent retraining. For daily data, forecasting with machine learning models usually costs less than with deep learning models. On average, the continuous retraining scenario costs approximately \$750,000 and this cost drops down to almost \$250,000 in the no retraining scenario, implying direct savings of more than 60%. Moreover, machine learning models allow for higher savings. Indeed, even if the models have to be updated, going from retraining every week to retraining every month allows to get direct savings of almost 75%, while it can be less than 30% for deep learning models. This implies that machine learning methods are a marginally better choice to reduce the costs of forecasting when coupled with less frequent model retraining as the frequency of the data increases. For weekly data, these differences are less evident, meaning that machine learning and deep learning models usually have very similar costs and savings profiles (without considering the Random Forest model, which is usually more than 10 times slower than other methods). This is to be expected since lower frequencies also imply smaller datasets and lower computing time. In this case, the average cost of forecasting under a continuous retraining scenario is \$250,000 (1/3 of that of daily data), and it drops to \$15,000 in the least frequent retraining scenario, producing direct savings of 90%. Nevertheless, for weekly data too, moving from retraining a model every week ($r=1$) to just once a month ($r=4$) allows to reduce the cost of almost 75%.
It may be argued that for a large retailer (like Walmart), these costs (and savings) may be negligible. However, it has to be noted that the cost reduction obtained from less frequent retraining comes with no shortcomings in terms of forecasting accuracy (especially for point forecasting). Indeed, as already mentioned, models under a periodic retraining scenario result in at least the same (if not even better) forecasting performance compared to the usual practice of continuous retraining. Moreover, periodic retraining may be a good practice even when probabilistic forecasts are needed, balancing costs and accuracy effectively.
Figure 6: Daily data estimated costs and percentage savings for each method and retrain scenario combination. The black line represents the average. Costs are in real values, while savings are expressed in relative terms with respect to the baseline scenario, $r=7$.
Figure 7: Weekly data estimated costs and percentage savings for each method and retrain scenario combination. The black line represents the average. Costs are in real values, while savings are expressed in relative terms with respect to the baseline scenario, $r=1$.
Finally, in this study, since we are dealing with global models, we assumed a fixed model retraining frequency across all forecast origins and time series. While our results showed no significant difference in forecasting accuracy when reducing the retraining frequency, we believe there is potential to further improve accuracy, especially when adopting some form of periodic retraining, for instance, by allowing the retraining frequency to vary dynamically. Such adaptive strategies could involve automated procedures that monitor forecast error levels and trigger updates only when errors exceed a certain threshold. This approach would still yield computational savings, albeit smaller, since the model re-training would not be required at every review point. In this direction, Figures 8 and 9 illustrate the ”optimal” retraining frequency identified for each series in the M5 and VN1 datasets, considering all the models, based on our experiments. The results confirm that although certain series may benefit from frequent retraining, in most cases, a less frequent retraining of the models leads to better accuracy and efficiency. Indeed, for point forecasting, the optimal retraining frequency was found to be between 3 and 4 weeks (approximately 21 observations) for the M5 dataset, and around 8 to 10 weeks for the VN1 dataset. In the context of probabilistic forecasting, instead, the optimal retraining intervals were shorter (around 2 weeks for M5 and 4 weeks for VN1), suggesting that uncertainty estimation may require more frequent updates to maintain accurate levels. Moreover, while these results do not differentiate among the global forecasting methods used in the experiment, individual model differences exist, implying that even further improvements may be obtained by relying on specific model (see Supplementary material). These findings suggest that exploring intermediate retraining strategies and determining the optimal retraining frequency offers promising avenues for reducing the computational cost of the forecasting process while maintaining a high level of accuracy.
Figure 8: Overall optimal retraining frequency based on RMSSE. For each series, the optimal frequency is identified based on the average accuracy reported across the evaluation rounds per update frequency, considering all the models.
Figure 9: Overall optimal retraining frequency based on SMQL. For each series, the optimal frequency is identified based on the average accuracy reported across the evaluation rounds per update frequency, considering all the models.
5 Conclusions
In this study, we went beyond the traditional evaluation of forecasting models by also exploring the computational cost associated with generating forecasts. We examined the effects of retraining frequency on the accuracy and computational efficiency of global time series forecasting models. By systematically evaluating various retraining scenarios, ranging from continuous retraining to no retraining at all, across multiple machine learning and deep learning models, we aimed to address a key question in forecasting: is frequent retraining necessary for maintaining the high predictive performance of global models? Our findings, derived from an extensive evaluation of ten different global models across two large-scale, real-world retail datasets, challenge the conventional wisdom of continuous model retraining. Indeed, our results indicate that the common practice of continuous retraining may not always be justified, and that periodic retraining strategies can offer substantial benefits in terms of computational efficiency without a significant loss in accuracy.
Our empirical evaluation provides evidence that the forecasting accuracy of global models remains stable even when the retraining frequency is significantly reduced. For point forecasting, the root mean squared scaled error (RMSSE) results indicate that models trained less frequently do not exhibit substantial performance degradation. In fact, in some cases, particularly for periodic retraining frequencies, we observe marginal improvements in accuracy. For probabilistic forecasting, the results are slightly more nuanced. While less frequent retraining does lead to minor reductions in accuracy as measured by the scaled Multi-Quantile Loss (SMQL), the degradation is relatively small (typically within 5-6%). This implies that for most practical applications, especially those where forecasting costs are key considerations, periodic retraining provides a favorable balance between accuracy and efficiency. This suggests that global models, that learn shared dynamics across multiple time series, are robust to the absence of frequent updates, especially in contexts with relatively stable demand patterns.
Considering periodic retraining scenarios allows for effective management of the computing time, and in turn, the costs of forecasting. Indeed, we have shown how the computational time can be directly translated into actual costs for the company. One of the most significant findings of this study is the exponential reduction in computational costs as retraining frequency decreases. The computational time (CT) analysis demonstrates that moving from continuous retraining to a monthly retraining schedule can reduce computational costs by approximately 75%. In the extreme case of no retraining, the cost reductions approach 90%, representing a resource-saving opportunity with no shortcomings. Furthermore, the cost analysis performed on realistic retail settings (e.g., Walmart-sized operations) reveals that periodic retraining can yield relevant financial savings. The estimated costs of forecasting, which scale with computing service expenses, decrease sharply when retraining is performed less frequently. This suggests that organizations can achieve economic benefits by optimizing their retraining schedules without compromising the forecast quality.
The implications of these findings are highly relevant for both academic research and industry applications. First, they challenge the prevailing assumption that forecasting models require frequent updates to maintain high predictive performance. Instead, our results suggest that global models remain effective over extended periods, meaning that retraining can be strategically planned rather than performed continuously. For practitioners, our results provide concrete guidelines on how often retraining should occur. In general, retraining every month appears to be a viable option to balance probabilistic accuracy and costs. Instead, if the forecasting objective is on point forecast, then even longer retraining scenarios may be adopted. Moreover, our results shed some light on the computational comparison between machine learning and deep learning models. We found that the former benefits more from less frequent retraining as the frequency of the data increases. This implies that, for large-scale applications, like the retail industry, where the forecasts of many different SKUs have to be provided regularly, machine learning models are a marginally better choice to reduce the costs of forecasting when coupled with less frequent model retraining strategies as the frequency of the data increases. These insights also have broader implications for sustainability in machine learning and AI-driven forecasting. The computational cost savings associated with reduced retraining directly translate into lower energy consumption, making forecasting operations more environmentally sustainable. This aligns with recent discussions on ”Green AI” which emphasizes the importance of optimizing computational resources to reduce the environmental impact of machine learning applications.
While our findings provide strong evidence for the feasibility of less frequent retraining in global forecasting models, some limitations remain. First, our study focused on two large retail datasets (M5 and VN1), which may not fully capture all possible time series patterns. Moreover, in this paper, we assumed that the data-generating process remains stable without exhibiting significant trends or concept drifts, but in some real-world applications, this assumption may simply not be valid. Future research could extend this analysis to other domains, such as financial time series or industrial production data, to verify the general validity of these findings. Furthermore, our study did not explore adaptive retraining strategies, where models are updated only when a significant drift in data distribution is detected. Such adaptive approaches could provide an even more refined balance between accuracy and computational efficiency. Future work could investigate the integration of drift detection mechanisms to optimize retraining schedules dynamically. Another direction for future studies may be the analysis of the effect of different retraining scenarios on the stability of global models, that is, answering the question ”Do global models produce stable forecasts even if not continuously retrained?”. Extending the analysis to some very recent global forecasting methods that demonstrate state-of-the-art performance (e.g., transformers) would also be a natural path for future research. Finally, we encourage further studies to explore the interaction between retraining frequency and hyperparameter optimization. Indeed, different models may respond differently to retraining strategies depending on their hyperparameter configurations, and understanding these interactions could lead to even more efficient forecasting strategies.
In summary, our study shows that frequent retraining is not necessarily required to maintain high forecasting accuracy in global models. Less frequent retraining can significantly reduce computational costs while maintaining competitive forecasting performance. These findings have important implications for organizations seeking to optimize their forecasting pipelines, offering a pathway toward more efficient and sustainable forecasting practices. By shifting from a continuous retraining paradigm to a periodic retraining approach, businesses can achieve relevant cost savings while still ensuring high predictive performance.
Supplementary material
In this section, we provide tables and figures related to the empirical results of the M5 and VN1 datasets. Note that results are obtained by averaging over the whole dataset after trimming almost 0.5% of the time series at the extremes of the distribution of each metric. This is a standard practice, since those time series that exhibit completely out-of-scale values of the evaluation metric should be modeled separately.
The Tables 3 and 4 show the forecast accuracy of the different models along the examined retrain scenarios for the M5 daily dataset, while Table 5 depicts the computing time in seconds.
The Tables in 6 and 7 show the forecast accuracy of the different models along the examined retrain scenarios for the VN1 weekly dataset, while 8 depicts the computing time in seconds.
Figures 10 and 11 show the results of the Friedman-Nemenyi test in the context of both point and probabilistic forecasting for the VN1 dataset.
The costs tables 9 and 10 show the estimated cost in real values of each scenario for daily and weekly data, respectively.
Figures 12 and 13 display the optimal retraining frequency for each model in the context of the M5 dataset for point and probabilistic forecasting. Figures 14 and 15 display the optimal retraining frequency for each model in the context of the VN1 dataset for point and probabilistic forecasting. For each series, the optimal frequency is identified based on the average accuracy reported across the evaluation rounds per retraining period.
Method 7 14 21 30 60 90 120 150 180 364
LR 0.760 0.760 0.760 0.759 0.759 0.759 0.759 0.759 0.760 0.759
XGBoost 0.738 0.738 0.738 0.738 0.738 0.738 0.737 0.737 0.737 0.736
LGBM 0.754 0.754 0.754 0.754 0.754 0.754 0.754 0.754 0.754 0.754
CatBoost 0.931 0.945 0.937 0.944 0.936 0.952 0.941 0.946 0.950 0.941
MLP 0.804 0.801 0.801 0.802 0.801 0.803 0.800 0.802 0.802 0.802
TCN 0.849 0.849 0.848 0.848 0.848 0.848 0.848 0.848 0.847 0.845
NBEATSx 0.798 0.797 0.797 0.807 0.798 0.798 0.797 0.798 0.798 0.797
NHITS 0.811 0.809 0.818 0.809 0.809 0.807 0.808 0.808 0.808 0.807
Table 3: M5 RMSSE values for each method and retrain scenario combination.
Method 7 14 21 30 60 90 120 150 180 364
LR 0.268 0.268 0.269 0.269 0.271 0.272 0.273 0.272 0.273 0.274
XGBoost 0.245 0.246 0.247 0.247 0.249 0.251 0.252 0.251 0.252 0.254
LGBM 0.247 0.247 0.248 0.248 0.249 0.250 0.250 0.250 0.250 0.252
CatBoost 0.282 0.273 0.283 0.276 0.278 0.274 0.270 0.276 0.273 0.272
MLP 0.263 0.264 0.265 0.266 0.269 0.272 0.272 0.272 0.275 0.277
TCN 0.272 0.273 0.274 0.275 0.277 0.279 0.280 0.280 0.282 0.284
NBEATSx 0.262 0.263 0.264 0.269 0.268 0.271 0.271 0.271 0.274 0.276
NHITS 0.267 0.266 0.274 0.268 0.271 0.273 0.274 0.274 0.277 0.279
Table 4: M5 SMQL values for each method and retrain scenario combination.
Method 7 14 21 30 60 90 120 150 180 364
LR 11,372 6,068 4,388 3,275 2,012 1,517 1,410 1,358 1,133 964
XGBoost 15,417 8,189 5,843 4,413 2,567 2,022 1,748 1,750 1,402 1,160
LGBM 44,428 23,835 17,279 13,245 8,332 6,656 5,802 5,966 4,971 4,256
CatBoost 10,423 5,521 3,907 2,901 1,693 1,285 1,082 1,113 891 709
MLP 17,583 14,044 12,864 12,219 11,411 11,156 11,080 10,972 10,711 10,639
TCN 33,363 27,960 25,931 25,102 23,721 23,332 23,046 23,033 22,825 22,591
NBEATSx 21,226 16,331 14,965 13,583 12,387 11,969 11,670 11,681 11,458 11,265
NHITS 21,969 16,684 14,863 13,682 12,390 11,929 11,722 11,712 11,440 11,219
Table 5: M5 CT values for each method and retrain scenario combination.
Method 1 2 3 4 6 8 10 13 26 52
LR 7.564 7.540 7.374 7.444 7.339 7.550 7.552 7.364 7.495 7.560
RF 1.871 1.873 1.872 1.868 1.867 1.866 1.861 1.863 1.852 1.840
XGBoost 1.898 1.871 1.884 1.865 1.842 1.851 1.809 1.833 1.851 1.797
LGBM 3.540 3.520 3.471 3.282 3.414 3.743 3.447 3.355 3.382 3.561
CatBoost 6.384 6.348 6.388 5.969 5.738 6.051 5.072 5.300 5.785 6.156
MLP 1.554 1.483 1.477 1.472 1.472 1.475 1.480 1.480 1.467 1.455
LSTM 1.960 1.960 1.960 1.960 1.960 1.960 1.960 1.960 1.960 1.960
TCN 1.960 1.960 1.960 1.960 1.960 1.960 1.960 1.960 1.960 1.960
NBEATSx 1.704 1.696 1.661 1.561 1.519 1.525 1.544 1.462 1.455 1.456
NHITS 1.705 1.680 1.629 1.556 1.547 1.460 1.528 1.453 1.455 1.451
Table 6: VN1 RMSSE values for each method and retrain scenario combination.
Method 1 2 3 4 6 8 10 13 26 52
LR 4.905 4.923 4.802 4.901 4.871 4.911 4.941 4.830 4.752 5.072
RF 1.220 1.230 1.232 1.224 1.237 1.254 1.257 1.284 1.313 1.326
XGBoost 1.503 1.496 1.493 1.505 1.483 1.488 1.466 1.451 1.515 1.479
LGBM 2.318 2.292 2.230 2.157 2.230 2.322 2.266 2.366 2.226 2.275
CatBoost 4.945 4.867 5.084 4.771 4.476 4.582 4.081 4.133 4.351 5.023
MLP 0.806 0.800 0.796 0.795 0.803 0.805 0.807 0.819 0.824 0.847
LSTM 0.892 0.895 0.899 0.903 0.910 0.917 0.924 0.935 0.957 1.003
TCN 0.891 0.895 0.899 0.903 0.910 0.917 0.924 0.935 0.957 1.003
NBEATSx 0.991 0.995 0.966 0.897 0.861 0.866 0.893 0.825 0.839 0.875
NHITS 0.992 0.996 0.944 0.894 0.876 0.814 0.880 0.820 0.833 0.862
Table 7: VN1 SMQL values for each method and retrain scenario combination.
Method 1 2 3 4 6 8 10 13 26 52
LR 236 121 89 66 50 39 33 34 23 17
RF 24,862 12,380 8,715 6,177 4,355 3,084 2,461 2,570 1,262 602
XGBoost 530 274 198 146 108 81 69 70 44 31
LGBM 7,824 4,037 2,894 2,144 1,576 1,188 1,007 1,027 630 429
CatBoost 805 412 287 207 149 109 90 93 51 31
MLP 962 534 409 327 265 223 202 202 159 138
LSTM 1,284 759 602 495 418 363 338 340 285 258
TCN 1,127 639 494 394 322 272 247 253 200 175
NBEATSx 1,244 694 531 421 342 285 258 259 203 175
NHITS 1,251 704 537 427 346 288 261 261 208 180
Table 8: VN1 CT values for each method and retrain scenario combination.
Figure 10: VN1 Friedman-Nemenyi test results based on RMSSE.
Figure 11: VN1 Friedman-Nemenyi test results based on SMQL.
Method 7 14 21 30 60 90 120 150 180 364 LR 390,732 208,499 150,791 112,547 69,132 52,131 48,474 46,678 38,937 33,151 XGBoost 529,679 281,358 200,778 151,635 88,200 69,495 60,082 60,155 48,189 39,861 LGBM 1,526,424 818,569 593,676 455,075 286,262 228,698 199,337 204,972 170,802 146,252 CatBoost 358,123 189,714 134,258 99,699 58,167 44,164 37,206 38,266 30,622 24,362 MLP 604,120 482,505 441,981 419,823 392,046 383,287 380,697 376,968 368,023 365,525 TCN 1,146,256 960,626 890,924 862,432 815,006 801,612 791,801 791,348 784,201 776,181 NBEATSx 729,263 561,105 514,166 466,683 425,593 411,232 400,953 401,352 393,673 387,058 NHITS 754,783 573,233 510,655 470,077 425,684 409,844 402,731 402,394 393,071 385,470 Average 754,922 509,451 429,654 379,746 320,011 300,058 290,160 290,267 278,440 269,733
Table 9: M5 estimated costs (in \$) for each method and retrain scenario combination.
Method 1 2 3 4 6 8 10 13 26 52 LR 15,234 7,839 5,731 4,292 3,205 2,508 2,140 2,190 1,463 1,099 RF 1,605,768 799,597 562,896 398,954 281,246 199,214 158,959 165,975 81,529 38,865 XGBoost 34,254 17,687 12,796 9,413 6,944 5,262 4,473 4,534 2,863 2,005 LGBM 505,304 260,721 186,904 138,460 101,775 76,738 65,013 66,335 40,721 27,701 CatBoost 52,021 26,594 18,515 13,366 9,611 7,022 5,798 5,977 3,311 2,023 MLP 62,102 34,489 26,431 21,149 17,121 14,391 13,027 13,015 10,255 8,891 LSTM 82,927 49,019 38,851 31,962 27,022 23,465 21,829 21,932 18,415 16,652 TCN 72,763 41,291 31,910 25,464 20,791 17,579 15,983 16,365 12,936 11,284 NBEATSx 80,337 44,837 34,283 27,179 22,057 18,387 16,686 16,702 13,085 11,277 NHITS 80,776 45,458 34,688 27,585 22,336 18,607 16,888 16,855 13,405 11,601 Average 259,149 132,753 95,301 69,782 51,211 38,317 32,080 32,988 19,798 13,140
Table 10: VN1 estimated costs (in \$) for each method and retrain scenario combination.
Figure 12: M5 optimal retraining frequency results for each model based on RMSSE.
Figure 13: M5 optimal retraining frequency results for each model based on SMQL.
Figure 14: VN1 optimal retraining frequency results for each model based on RMSSE.
Figure 15: VN1 optimal retraining frequency results for each model based on SMQL.